# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v4

@@ 1 Introduction
The rest of this paper is organized as follows. Section 2 reviews state-of-the-art physics-aware ML. Section 3 presents process-based hydrology in physics-aware ML (hydroPML), and its applications in real-time flood forecasts. Section 4 concludes the paper.
@@ 2 State-of-the-art Physics-aware Machine Learning
Physics-aware ML (PaML) aims to take the best from both physics-based modeling and state-of-the-art ML models to better solve scientific problems. As shown in Fig. 2, a structured community of existing PaML methodologies that integrate prior physical knowledge or physics-based modeling into ML is built. We categorize PaML approaches into four groups based on the way physics and ML are combined, including physical data-guided ML (PDgML), physics-informed ML (PiML), physics-embedded ML (PeML), and physics-aware hybrid learning (PaHL). Each is discussed in detail below.
@@ 2.1 Physical Data-guided Machine Learning
Table 1: Proposed Physical Data-guided Machine Learning
Categories Types Examples Objective(s) Limitation(s)
Informer [33] Long sequence time-series forecasting
ConvLSTM [34] A spatio-temporal sequence forecasting problem
Geometric deep learning [36] A spatially-irregular forecast method
Neural message passing-based models [39, 40, 41, 42, 43, 44, 45] A general framework for supervised learning on graphs (arbitrary domains) Mesh-based or particle-based; message-passing between small-scale moving and interacting objects
Fourier neural operator (FNOs) [50] Geo-FNO [51], Domain agnostic FNO [52] Solve PDEs on arbitrary geometries Unstable, limited generalization, the basis functions in FFTs are generally frequency localised with no spatial resolution
Physics-discovery Neural Networks Physics-discovery with unavailable prior knowledge Symbolic neural networks (SNNs) [58, 59, 60] Infer concise equations without any prior knowledge Challenging task due to complex and nonlinear systems, as well as noisy and incomplete data
AI Feynman [61, 62] Discover generalized symmetries in complex data
GNN [63] Identify the nontrivial relationship due to its exceptional inductive capability
PDE-FIND [65] Data-driven discovery for PDEs
PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data. PDgML consists of one or a combination of deep neural networks (DNN) [66], convolutional neural networks (CNNs) [67], recurrent neural networks (RNNs) [68], graph neural networks (GNN) [69], generative adversarial networks (GAN) [70], Transfomer [71], deep reinforcement learning (DRL) [72], deep operator networks [73], and physics-discovery neural networks [74]. The objectives and limitations of different physical data-guided ML are briefly discussed in Table 1.
@@ 2.1.1 Classical Physical Data-guided Neural Networks
Spatio-temporal modeling. DL has become a powerful tool for solving complex problems involving large datasets. Today, most of the data generated by smart devices are time series, and predicting these data is a common issue in almost every research field [75]. The Elman network, recognized as the pioneering RNN, introduce the concept of incorporating the $t$ state of a hidden unit to enhance sequence prediction [76]. To address the vanishing gradient issue inherent in RNNs, long short-term memory (LSTM) networks are developed [31]. These LSTM networks utilize three distinct gates to retain crucial information over extended periods while discarding what is deemed irrelevant. Physics-guided LSTM [32] is developed to handle extreme events and monotonic relationships in rainfall-runoff simulations using deep learning. It incorporates three physical mechanisms to regulate the LSTM network, ensuring it effectively manages high flow, low flow, and monotonic properties. However, current methods are designed within restricted problem settings, such as predicting fewer than 48 points [77]. Recent studies highlight the potential of Transformer models to improve prediction capabilities. In response, Informer [33] has been introduced to address long sequence time-series forecasting, demonstrating the Transformer-like model’s effectiveness in capturing long-range dependencies between inputs and outputs in lengthy sequences.
@@
Additionally, an army of physical data-guided spatio-temporal models [78, 79, 80] is proposed by imposing temporal and spatial inductive biases based on RNNs and CNNs. For example,ConvLSTM [34] is a recurrent neural network designed for spatio-temporal prediction, incorporating convolutional layers in both the input-to-state and state-to-state transitions. Furthermore, inspired by the strong representation capacity of Transformer architectures (such as bidirectional encoder representations from transformers (BERT) [81] in natural language processing, vision Transformer in computer vision [82, 83]), quite a few physics-guided Transformer-based frameworks are presented. For example, a space-time Transformer for Earth system forecasting, Earthformer [35], is presented based on an efficient space-time attention block. A global data-driven weather forecasting model, FourCastNet [84], is proposed to provide accurate short- to medium-range global predictions at 0.25°$\degree$ resolution. In addition, in order to improve prediction accuracy and uncertainty, irregular space-time grid prediction problems, and other issues, a collection of classic spatio-temporal models are proposed, such as GraphCast [85] based on GNN for medium-range global weather forecasting, physically constrained GANs for improving local distributions and spatial structure of precipitation fields [86], and spatially-irregular forecasting based on geometric deep learning [36].
@@
Dynamic system modeling. Simulating multi-dimensional PDE systems with data-driven NNs has been a renewed research topic, that dates back to the last century [87]. PDgML for dynamic system modeling aims to learn a PDE model (if available) from a provided dataset and make accurate and efficient predictions using the learned model. Development has taken place in roughly three phases.
@@
First, any early attempt [88] at the data-driven discovery of hidden physical laws compares numerical differentiations of the experimental data with analytic gradients of candidate functions. This approach employs symbolic regression and an evolutionary algorithm to determine the nonlinear dynamic system. Recently, motivated by the latest development of neural networks, a fully non-mechanistic method, PDE-Net [37], is designed to predict the dynamics of complex systems accurately and uncover the underlying hidden PDE models based on learning convolution kernels (filters). Further works have extended the approach with residual CNNs [89], symbolic neural networks [90], feed-forward networks [91], and deep fully convolutional networks [92]. However, these models are fundamentally constrained by their reliance on discretizing the input domain using a sample-inefficient grid. As a result, they struggle to effectively manage temporally or spatially sparse and non-uniform observations, which are frequently encountered in practical applications. Recently, some convolution operators, such as Mesh-Conv [38], have been proposed to model non-uniform structured or unstructured spatial mesh.
@@
Second, neural message passing-based models such as GNN-based model [39], autoregressive solver [40], and contrastive learning-based message passing graph neural networks (MP-GNNs) [41], are related to interaction networks where the state of an object evolves as a function of its neighboring objects, forming dynamic relational graphs instead of grids. Neural message passing is a general framework for supervised (physical data-guided) learning on graphs [42]. In addition to the common mesh-based neural message passing-based approaches, particle-based frameworks have attracted considerable interest in scientific modeling for physical systems, particularly those employing GNNs [43, 44, 45]. Although neural message passing-based models can achieve geometry-adaptive learning of nonlinear PDEs with arbitrary domains, these models can apply message-passing between small-scale moving and interacting objects.
@@
Third, DRL has been increasingly utilized in various data-driven dynamic system modeling due to its capability to address complex decision-making problems characterized by non-linearity and high dimensionality [93, 94]. For instance, a DRL-based method [46] has been developed to manage a coupled 2D system with both fluid and rigid bodies, utilizing a position-based reward function.
@@ 2.1.2 Deep Operator Networks
Operator learning strategies are based on the principles of the universal approximation theorem [47]. This capability to approximate functions forms the foundation of modern developments in deep operator networks. Here, we present an introduction to deep operator networks (DeepONet) [47], the Fourier neural operator (FNOs) [50], the wavelet neural operators (WNOs) [54], and the graph neural operators (GNOs) [56].
@@
FNOs [50, 98] use Fourier Transformations to move to the infinite-dimensional response space. These operators process input functions defined on a well-defined, equally spaced lattice grid and produce the desired fields at the same grid points. The network parameters are established and trained in the Fourier domain instead of the physical space. FNO has become one of the mainstream physical data-guided deep operator networks used to solve PDEs [73]. Specifically, to solve PDEs on arbitrary geometries, Geo-FNO [51] is designed by learning to deform the input (physical) domain. Domain agnostic FNO [52] is proposed by incorporating a smoothed characteristic function within the integral layer architecture of FNOs. However, the FNO and the geo-FNO perform worse and are unstable on complex geometries and noisy data. Thus, factorized FNO [53] is proposed to improve representation stability for the operator. In addition, a host of variants of FNOs are proposed to improve the generalization of FNOs for solving PDEs, such as incremental FNO [99].
@@
GNOs [56] utilize message passing on graph networks to capture the non-local structure of data. Graph kernel networks (GKNs), which originate from parameterizing Green’s functions in an iterative architecture [56], are primarily employed in GNOs. Similar to FNOs, GKNs consist of a lifting layer, iterative kernel integration layers, and a projection layer. However, recent studies have noted that GKNs may exhibit instability with an increased number of iterative kernel integration layers [57]. Therefore, a resolution-independent nonlocal neural operator [57] has been proposed to address these issues.
@@ 2.1.3 Physics-discovery Neural Networks
Discovering dynamical processes is crucial as it enables us to uncover the fundamental physical laws that govern complex systems. However, the discovery of the governing equations presents a significant challenge due to the inherent complexity and nonlinearity of these systems, as well as the presence of noisy and incomplete data. With the rapid advancement of machine intelligence [12], the past three decades have witnessed significant development in physics-discovery neural networks, at the intersection of ML and scientific discovery. Two main approaches have emerged based on unavailable and available prior knowledge of the physics system. First, the natural and optimal solution to automatically identifying the governing equations for dynamic systems is to learn a symbolic model from experimental data. Symbolic regression [101] and symbolic neural networks (SNNs) [58, 59, 60] have been explored extensively for inferring concise equations without prior knowledge. A physics-inspired method called AI Feynman [61, 62] is proposed for symbolic regression. AI Feynman employs NNs to discover generalized symmetries in complex data, enabling the recursive decomposition of difficult problems into simpler ones with fewer variables. GNN [63] is also utilized to identify the nontrivial relation due to its exceptional inductive capability. In addition, PDgML is applied for knowledge discovery that cannot be modeled from a physical process-based perspective, such as the two-way feedback between human and water systems [6, 7]. Second, discovering dynamic behavior from data often involves defining a large set of possible mathematical basis functions or process-based models based on prior knowledge. Representative works include SINDy [64] and PDE-FIND [65] for ordinary differential equations (ODEs) and PDEs, respectively. However, these standard sparse representation-based methods are typically limited to high-fidelity noiseless measurements, which are often difficult and expensive to obtain. Consequently, recent efforts have focused on discovering PDEs from sparse or noisy data [102, 103].
@@ 2.2 Physics-informed Machine Learning
In scientific computing, physical phenomena are typically described using a robust mathematical framework that includes governing differential equations, as well as initial and boundary conditions. PiML is a widely-used approach that integrates physical laws into ML models (for example, designing loss functions or regularization), facilitates the accurate capture of dynamic patterns and concomitantly diminishes the search space for model parameters. This approach is sometimes referred to as imposing differentiable constraints in loss functions. It integrates (noisy) data and mathematical models, and implements them through neural networks (physics-informed neural networks) or kernel-based neural operators (physics-informed neural operators). The objectives and limitations of different physics-informed machine learning are briefly discussed in Table 2.2.
Table 2: Proposed Physics-informed Machine Learning
Categories Types Examples Objective(s) Limitation(s)
CINN [109] A straightforward and efficient characteristic-informed neural networks for solving forward and inverse problems in hyperbolic PDEs Limited application; limited expression and generalization ability
PI-PINN [121] Orthogonal polynomials are used to construct the neural network for solving nonlinear partial differential equations Limited expressivity of framework
Improving speed and efficiency Meta-learning based PINNs de Avila Belbute-Peres et al. [125] meta-trains a hyper-network that for each task can generate weights for a small neural network. Psaros et al. [126] meta-learns a loss function that is used to optimize the NN. Penwarden et al. [127] proposes a meta-learning approach that learns an initialization of weights such that the NN can be optimized quickly. Limited expressivity of framework, unsteady in solving related PDEs
PINN-SR Physics-informed neural network with sparse regression to discover governing partial differential equations from scarce and noisy data for nonlinear spatio-temporal systems High computational cost need a vast number of collocation points for complex systems; limitations of PINNs
Physics- informed Neural Operators WNO-based Physics-informed WNO [143] A physics-informed WNO for learning solution operators of families of parametric PDEs without labeled training data, leveraging the time-frequency localization advantage of wavelets. Complexity; the absence of a standardized procedure to obtain the optimum and desired network architecture of a physics-informed WNO and to assign the weights of boundary/initial condition loss; the lack of information regarding how to learn the operator for systems involving multi-scale physics systems

### v5

@@ 1 Introduction
The rest of this paper is organized as follows. Section 2 reviews state-of-the-art physics-aware ML. Section 3 presents process-based hydrology in physics-aware ML (HydroPML), and its application highlights. Section 4 concludes the paper.
@@ 2 State-of-the-art Physics-aware Machine Learning
PaML aims to integrate the strengths of physics-based modeling and state-of-the-art ML models to effectively tackle scientific challenges. As shown in Fig. 2, a structured community of existing PaML methodologies that integrate prior physical knowledge or physics-based modeling into ML is built. We categorize PaML approaches into four groups based on the way physics and ML are combined, including physical data-guided ML (PDgML), physics-informed ML (PiML), physics-embedded ML (PeML), and physics-aware hybrid learning (PaHL). These four methods in the PaML community, including their corresponding benefits and drawbacks for scientific problems, are summarized in Table 1. Each is discussed in detail below.
Table 1: Four approaches to physics-aware machine learning.
Term Description Methods Links to physics Benefits and drawbacks
Physical Data-guided Machine Learning (PDgML) PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw training data. Classical physical data-guided neural networks (spatio-temporal modeling, dynamic system modeling), deep operator networks, physics-discovery neural networks. Learning from physical data (time series data of physical processes, dynamic process data, and physical process observations). [Benefits] Only need simulations from physics model, fast inference, the discovery of the dynamic behavior;
Physics-informed Machine Learning (PiML) PiML is a widely used approach to incorporating physical constraints that can be trained from additional information obtained by enforcing the physical laws (for example, designing loss functions or regularization). Physics-informed neural networks (PINNs), physics-informed deep neural operators (PINOs). Integrating physical laws (e.g., parametrized PDEs) by designing loss functions or regularization (e.g., PDE residual loss). [Benefits] Incomplete models and imperfect data, strong generalization in the limited data, understanding DL by the optimization process, tackling high dimensionality; uncertainty quantification, data-physics-driven parameters discovery;
Physics-embedded Machine Learning (PeML) PeML is achieved by embedding physics in the model frameworks or modules. Physical equation-embedded machine learning (PEeML), physical property-embedded machine learning (PPeML), physical condition-embedded machine learning (PCeML). Embedding physics, including differential equations, physical properties (e.g., conservation, symmetry, causality), and physical conditions (e.g., boundary and initial conditions in PDEs, boundaries of physical processes) into ML frameworks or modules. [Benefits] Interpretability of architecture and learning process, generalization;
Physics-aware Hybrid Learning (PaHL) PaHL directly combines pure physics-based models, such as numerical methods, climate, land, hydrology, and Earth system models with ML models. Depending on the hybrid approach, hybrid learning can be categorized as serial, parallel, or complex. Neural–numerical hybrid learning, neural–hydrology hybrid learning. Hybrid learning by combining physics-driven models and data-driven models. [Benefits] Can leverage both the flexibility of data-driven models along with the interpretability and generalizability of physics-driven models, speed, and operational convenience;
@@ 2.1 Physical Data-guided Machine Learning
PDgML is a supervised DL model that statistically learns the known or unknown physics of a desired phenomenon by extracting features or attributes from raw physical data. These physical data typically encompass three aspects. First, there are time series data obtained from physical process models, including catchment or global scale rainfall-runoff time series data from datasets such as catchment attributes and meteorology for large-sample studies (CAMELS) [31] and a global community dataset for large-sample hydrology (Caravan) [32], among others, as well as climate and weather reanalysis data and products like ERA5 [33]. Second, these physical data include dynamic process data derived from physical dynamic models, such as various PDE numerical solution data like PDEBench [34] and hydrodynamic data from flood dynamics simulations [35]. Finally, these physical data comprise other relevant physical observations for hydrological processes and events, such as remote sensing data, hydrological observations, and measurement data for floods and landslides [36]. Furthermore, PDgML consists of one or a combination of deep neural networks (DNN) [37], convolutional neural networks (CNNs) [38], recurrent neural networks (RNNs) [39], graph neural networks (GNN) [40], generative adversarial networks (GAN) [41], Transfomer [42], deep reinforcement learning (DRL) [43], deep operator networks [44], and physics-discovery neural networks [45]. The objectives and limitations of different physical data-guided ML are briefly discussed in Supplementary Table 1 of Supplementary Material (SM).
@@ 2.1.1 Classical Physical Data-guided Neural Networks
Spatio-temporal modeling. DL has become a powerful tool for solving complex problems involving large physical data. For the time series data obtained from physical process models, such as rainfall-runoff data, a series of time series modeling networks are proposed to extract features through supervised training. To address the vanishing gradient issue inherent in RNNs, long short-term memory (LSTM) networks are developed [46]. These LSTM networks utilize three distinct gates to retain crucial information over extended periods while discarding what is deemed irrelevant. For modeling storage effects in hydrological processes, the advantage of LSTM networks is essential due to their ability to learn long-term dependencies between network inputs and outputs [10]. Physics-guided LSTM [47] is developed to handle extreme events and monotonic relationships in rainfall-runoff simulations using DL. It incorporates three physical mechanisms to regulate the LSTM network, ensuring it effectively manages high flow, low flow, and monotonic properties. However, current methods are designed within restricted problem settings, such as predicting fewer than 48 points [48]. Recent studies highlight the potential of Transformer models to improve prediction capabilities. In response, Informer [49] has been introduced to address long sequence time-series forecasting, demonstrating the Transformer-like model’s effectiveness in capturing long-range dependencies between inputs and outputs in lengthy sequences.
@@
Additionally, an army of physical data-guided spatio-temporal models [50, 51, 52] is proposed by imposing temporal and spatial inductive biases based on RNNs and CNNs. These physical data-guided spatio-temporal models integrate spatio-temporal data from physical models or observations, such as climate and weather reanalysis data and hydrological observations, with data-driven spatio-temporal models. These models achieve robust and accurate spatio-temporal predictions of physical variables, through effective supervised training. For example, ConvLSTM [53] is a recurrent neural network designed for spatio-temporal prediction, incorporating convolutional layers in both the input-to-state and state-to-state transitions. DisasterNets [54] is a type of CNN for spatio-temporal disaster mapping based on Earth observation data. Furthermore, inspired by the strong representation capacity of Transformer architectures (such as bidirectional encoder representations from transformers (BERT) [55] in natural language processing, vision Transformer in computer vision [56, 57]), quite a few physics-guided Transformer-based frameworks are presented based on physical simulation data. For example, a space-time Transformer for Earth system forecasting, Earthformer [58], is presented based on an efficient space-time attention block. A global data-driven weather forecasting model, FourCastNet [59], is proposed to provide accurate short- to medium-range global predictions at 0.25°$\degree$ resolution. In addition, in order to improve prediction accuracy and uncertainty, irregular space-time grid prediction problems, and other issues, a collection of classic spatio-temporal models are proposed, such as GraphCast [60] based on GNN for medium-range global weather forecasting, physically constrained GANs for improving local distributions and spatial structure of precipitation fields [61], and spatially-irregular forecasting based on geometric deep learning [62].
@@
Dynamic system modeling. Simulating multi-dimensional PDE systems with data-driven NNs has been a renewed research topic, that dates back to the last century [63]. PDgML for dynamic system modeling aims to learn a PDE solution (if available) from a provided simulation dataset and make accurate and efficient predictions using the learned model. Development has taken place in roughly three phases.
@@
First, any early attempt [64] at the data-driven discovery of hidden physical laws compares numerical differentiations of the experimental data with analytic gradients of candidate functions. This approach employs symbolic regression and an evolutionary algorithm to determine the nonlinear dynamic system. Recently, motivated by the latest development of NNs, a fully non-mechanistic method, PDE-Net [65], is designed to predict the dynamics of complex systems accurately and uncover the underlying hidden PDE models based on learning convolution kernels (filters). Specifically, PDE-Net [65] utilizes the forward Euler method for temporal discretization, where each time step iteration is approximated to a nonlinear function using convolution operators with appropriately constrained filters. The nonlinear PDE system is then iteratively solved by stacking multiple convolution operators with shared weights. Further works have extended the approach by employing different networks to approximate the nonlinear function, including residual CNNs [66], symbolic neural networks [67], feed-forward networks [68], and deep fully convolutional networks [69]. However, these models are fundamentally constrained by their reliance on discretizing the input domain using a sample-inefficient grid. As a result, they struggle to effectively manage temporally or spatially sparse and non-uniform observations, which are frequently encountered in practical applications. Recently, some convolution operators, such as Mesh-Conv [70], have been proposed to model non-uniform structured or unstructured spatial mesh. Mesh-Conv [70] is achieved by introducing local weights to decompose standard convolutions, thereby integrating data structure information of spatial mesh.
@@
Second, neural message passing-based models such as GNN-based model [71], contrastive learning-based message passing graph neural networks (MP-GNNs) [72], and autoregressive solver [73], are related to interaction networks where the state of an object evolves as a function of its neighboring objects, forming dynamic relational graphs instead of grids. Neural message passing is a general framework for supervised (physical data-guided) learning on graphs [74, 73]. We illustrate the PDE solution process using the message passing neural PDE solver [73] as an example. The discrete physical grid is constructed as a graph with nodes and edges, where nodes represent grid cells and edges define local neighborhoods. Employing the encode-process-decode framework for autoregressive solver, the encoder computes node embeddings, and the processor executes multiple time steps of learned message passing. Subsequently, the decoder generates predictions for the next timestep across spatial locations on the physical grid. In addition to the common mesh-based neural message passing-based approaches, particle-based frameworks have attracted considerable interest in scientific modeling for physical systems, particularly those employing GNNs [75, 76, 77]. Although neural message passing-based models can achieve geometry-adaptive learning of nonlinear PDEs with arbitrary domains, these models can apply message-passing between small-scale moving and interacting objects.
@@
Third, DRL has been increasingly utilized in various data-driven dynamic system modeling due to its capability to address complex decision-making problems characterized by non-linearity and high dimensionality [78, 79]. For instance, a DRL-based method [80] has been developed to manage a coupled 2D system with both fluid and rigid bodies, utilizing a position-based reward function. Initially, the fluid’s velocity field is extracted using a convolutional autoencoder. Encoded velocity and other fixed features are subsequently fed into a multilayer perceptron (MLP) to determine actions. Both the MLP and autoencoder are trained using the DRL-based method [80] to achieve physically realistic animations.
@@ 2.1.2 Deep Operator Networks
Deep operator networks constitute an emerging class of PDgML algorithms. Unlike traditional NNs that approximate solutions directly, these networks approximate mathematical operators, enabling faster solvers for dynamic processes represented by PDEs. Consequently, deep operator networks offer a robust alternative for extracting feature representations from dynamic process data, effectively addressing physical problems such as fluid dynamics equations and general PDE solutions. Operator learning strategies are based on the principles of the universal approximation theorem [81]. This capability to approximate functions forms the foundation of modern developments in deep operator networks. Here, we present an introduction to deep operator networks (DeepONet) [81], the Fourier neural operator (FNOs) [82], the wavelet neural operators (WNOs) [83], and the graph neural operators (GNOs) [84].
@@
FNOs [82] use Fourier Transformations to move to the infinite-dimensional response space. These operators process input functions defined on a well-defined, equally spaced lattice grid and produce the desired fields at the same grid points. The network parameters are established and trained in the Fourier domain instead of the physical space. FNO has become one of the mainstream physical data-guided deep operator networks used to solve PDEs [44]. Specifically, to solve PDEs on arbitrary geometries, Geo-FNO [90] is designed by learning to deform the input (physical) domain. Domain agnostic FNO [91] is proposed by incorporating a smoothed characteristic function within the integral layer architecture of FNOs. However, the FNO and the geo-FNO perform worse and are unstable on complex geometries and noisy data. Thus, a host of variants of FNOs (such as factorized FNO [92]) are proposed to improve the generalization and stability of FNOs for solving PDEs.
@@
GNOs [84] utilize message passing on graph networks to capture the non-local structure of data. Graph kernel networks (GKNs), which originate from parameterizing Green’s functions in an iterative architecture [84], are primarily employed in GNOs. Similar to FNOs, GKNs consist of a lifting layer, iterative kernel integration layers, and a projection layer. However, recent studies have noted that GKNs may exhibit instability with an increased number of iterative kernel integration layers. Therefore, a resolution-independent nonlocal neural operator [95] has been proposed to address these issues.
@@ 2.1.3 Physics-discovery Neural Networks
Discovering dynamical processes is crucial as it enables us to uncover the fundamental physical laws that govern complex systems. However, the discovery of the governing equations presents a significant challenge due to the inherent complexity and nonlinearity of these systems, as well as the presence of noisy and incomplete data. With the rapid advancement of machine intelligence [12], the past three decades have witnessed significant development in physics-discovery neural networks, at the intersection of ML and scientific discovery. Physics-discovery neural networks facilitate the connection between data-driven models and physics learning by discovering and learning underlying physical principles and dynamic behaviors directly from complex physical data.
Two main approaches have emerged based on unavailable and available prior knowledge of the physics system. First, the natural and optimal solution to automatically identifying the governing equations for dynamic systems is to learn a symbolic model from experimental data. Symbolic regression [96] and symbolic neural networks [97, 98] have been explored extensively for inferring concise equations without prior knowledge. A physics-inspired method called AI Feynman [99, 100] is proposed for symbolic regression. AI Feynman employs NNs to discover generalized symmetries in complex data, enabling the recursive decomposition of difficult problems into simpler ones with fewer variables. GNN [101] is also utilized to identify the nontrivial relation due to its exceptional inductive capability. In addition, PDgML is applied for knowledge discovery that cannot be modeled from a physical process-based perspective, such as the two-way feedback between human and water systems [6, 7]. Second, discovering dynamic behavior from data often involves defining a large set of possible mathematical basis functions or process-based models based on prior knowledge. Representative works include Sparse Identification of Nonlinear Dynamics (SINDy) [102] and PDE functional identification of nonlinear dynamics (PDE-FIND) [103] for ordinary differential equations (ODEs) and PDEs, respectively. However, these standard sparse representation-based methods are typically limited to high-fidelity noiseless measurements, which are often difficult and expensive to obtain. Consequently, recent efforts have focused on discovering PDEs from sparse or noisy data [104].
@@ 2.2 Physics-informed Machine Learning
In scientific computing, physical phenomena are typically described using a robust mathematical framework that includes governing differential equations, as well as initial and boundary conditions. PiML is a widely-used approach that integrates physical laws into ML models (for example, designing loss functions or regularization), facilitates the accurate capture of dynamic patterns and concomitantly diminishes the search space for model parameters. This approach is sometimes referred to as imposing differentiable constraints in loss functions. It integrates (noisy) data and mathematical models, and implements them through neural networks (physics-informed neural networks) or kernel-based neural operators (physics-informed neural operators). The objectives and limitations of different physics-informed machine learning are briefly discussed in Supplementary Table 2 of SM.